OpenStudy (anonymous):

If the monthly payment formula M= Pr(1+r)^n / (1+r)-1 ,what value would you put for "r" if the interest rate is 5.7%?

Nnesha (nnesha):

you have to change percent to decimal to get interest rate

OpenStudy (anonymous):

@Nnesha so itd be 0.0057?

Nnesha (nnesha):

how did you get 0.0057 ??

OpenStudy (anonymous):

@Nnesha i divided 5.7 over 100 isnt that how you do it?

Nnesha (nnesha):

okay so 5.7/100 = ??

OpenStudy (anonymous):

@Nnesha my calculator says 0.057 •.• what did you get?

OpenStudy (anonymous):

i got 0.057..... my calculator doesnt lie js! xD

Nnesha (nnesha):

and that's right

Nnesha (nnesha):

i got same but look at your first comment it's .057 not .0057 :)

OpenStudy (anonymous):

@Nnesha i have the choices between A) 0.00475 B) 0.0057 C) 0.475 D) 5.7

Nnesha (nnesha):

hmmmm r = interest which is 5.7% right but when you solve the equation to find M or P then you have to divide interest rate by 100 hmmmm

OpenStudy (anonymous):

@Nnesha yeaah so i just picked 0.0057 because its the closest

Nnesha (nnesha):

noo that's not the closest one if you divide 5.7 by 1000 then you would get .0057 hmmm let's see @jdoe0001 hmmm ^^^^^^^^^^^^^^

OpenStudy (jdoe0001):

heheh

OpenStudy (jdoe0001):

the "r" in the formula, uses a decimal form is a "per cent", or per hundreds thus you'd divie 5.7 by 100 to get the decimal form

Nnesha (nnesha):

http://prntscr.com/6rt6ct

jimthompson5910 (jim_thompson5910):

Annual Interest Rate = 5.7% Monthly Interest Rate = 5.7/12 = 0.475%

OpenStudy (anonymous):

@jim_thompson5910 robert purchased a 2010 model sedan for $16,000. The dealership offered him a $99/month payment for 60 months, after which the unpaid balance is due. If the interest rate is 6%, find the balloon payment due at the end of 60 months.

Nnesha (nnesha):

jim but n suppose to be 12 :( hmmmm ???